You are performing a right-tailed test with test statistic t=2.02t=2.02 and a sample size of 21. Find the p-value to 4 decimal places.

p-value:
Answer #1

Solution :

This is the right tailed test .

df = 21 - 1 = 20

P-value = P(t 20 > 2.02) = 0.0285